### v4

# Ramsey Number Counterexample Checking and One Vertex Extension Linearly Bound by $s$ and $t$

Abstract
The Ramsey number $R(s,t)$ is the smallest integer $n$ such that all graphs of size $n$ contain a clique of size $s$ or an independent set of size $t$. $\mathcal{R}(s,t,n)$ is the set of all counterexample graphs without this property for a given $n$. I prove that if a graph $G_{n+1}$ of size $n+1$ has $\max\{s,t\}+1$ subgraphs in $\mathcal{R}(s,t,n)$, then $G_{n+1}$ is in $\mathcal{R}(s,t,n+1)$. Based on this, I introduce algorithms for one-vertex extension and counterexample checking with runtime linearly bound by $s$ and $t$. The utility of these algorithms is shown by verifying $\mathcal{R}(4,6,36)$ and $\mathcal{R}(5,5,43)$ are empty given current sets $\mathcal{R}(4,6,35)$ and $\mathcal{R}(5,5,42)$.
1 Introduction
The Ramsey number $R(s,t)$ is defined as the smallest positive integer $n$ such that all graphs of size $n$ contain a clique of size $s$ or an independent set of size $t$. The field of Ramsey theory explores finding exact values, upper bounds and lower bounds for given $s$ and $t$ values. The field began with the proof that $R(s,t)$ must exist for all given $s$ and $t$, and has transformed into a rich and well-studied subset of combinatorics over the course of the last few decades [1, 2, 3]. An extensive survey with current bounds, applications, and directions of the field can be found at [4].
Lower bounds for a given $R(s,t,n)$ have been determined through direct mathematical proofs since the beginning of the field, and in the last 30 years saw success through the addition of computational approaches [4]. A counterexample graph for $R(s,t)\leqslant n$ is a graph of size $n$ such that there exist no cliques of size $s$ or independent sets of size $t$. $\mathcal{R}(s,t,n)$ is the set of all counterexample graphs to $R(s,t)$ for a given $n$.
Some basic observations are as follows: Denote $k:=|\mathcal{R}(s,t,n)|$. $k>0\Rightarrow R(s,t)>n$. $k=0$ and $|\mathcal{R}(s,t,n-1)|>0\Rightarrow R(s,t)=n$. The discovery of nonempty counterexample sets has been used extensively to improve lower bounds, with $R(4,6)>35$ and $R(5,5)>42$ being current examples [5, 6]. A full database of known counterexample graphs can be found at [7]. The sets of known counterexamples for $\mathcal{R}(5,5,42)$ and $\mathcal{R}(4,6,35)$ are not provably exhaustive, although McKay and Radziszowski conjecture that $\mathcal{R}(5,5,42)$ is [5].
McKay and Radziszowski explicitly state that the known 656 $\mathcal{R}(5,5,42)$ counterexamples cannot be used to generate any $\mathcal{R}(5,5,43)$ counterexamples [5]. While no exact algorithm is provided, it can be assumed they used a variation of the one-vertex extension algorithm used in [3]. Exoo did not explicitly mention the use of an algorithm proving no $\mathcal{R}(4,6,36)$ could be generated from the known 37 $\mathcal{R}(4,6,35)$ counterexamples [6].
The one-vertex extension algorithm developed in [3] has an exact description for $R(4,5)$ but would need modification for other $s$ and $t$ values. Additionally, the algorithm has a runtime bound by the number of triangles and independent sets of size 4 in each counterexample, meaning for a given counterexample the runtime is bound exponentially to $\max\{s,t\}-1$.
1.1 Contribution
I provide and prove the main theorem in Section 2. Using this, I first introduce a new algorithm for counterexample checking in Section 3. This algorithm has a runtime bound linearly by $\max\{s,t\}$. Then, in Section 4, I show a new algorithm for one-vertex extension, or going from $\mathcal{R}(s,t,n)$ to $\mathcal{R}(s,t,n+1)$. The algorithm has a runtime linearly bound by $\max\{s,t\}$ instead of exponentially. The algorithm also works independent of choice of $s$, $t$, and $n$. Given an input of a known $\mathcal{R}(s,t,n)$, it generates the set of all $\mathcal{R}(s,t,n+1)$ with at least 2 subgraphs within the known $\mathcal{R}(s,t,n)$, as opposed to prior algorithms that generate the set of all $\mathcal{R}(s,t,n+1)$ with at least 1 subgraph within the known $\mathcal{R}(s,t,n)$. This limits the utility of the algorithm in searching a broader space but allows for efficient execution. I validate this by verifying $\mathcal{R}(5,5,43)$ and $\mathcal{R}(4,6,37)$ are empty given current sets $\mathcal{R}(5,5,42)$ and $\mathcal{R}(4,6,36)$ on a single PC in under 15 minutes.
2 Main Theorem
Let $G_{n+1}$ be a simple, undirected graph of size $n+1$.
Theorem 1.
If $G_{n+1}$ has $\max\{s,t\}+1$ subgraphs in $\mathcal{R}(s,t,n)$, $G_{n+1}\in\mathcal{R}(s,t,n+1)$
Proof.
Assume $G_{n+1}$ has $\max\{s,t\}+1$ subgraphs in $\mathcal{R}(s,t,n)$. Each subgraph in $\mathcal{R}(s,t,n)$ must be of size $n$, and so exclude a vertex in $G_{n+1}$. Let $S=\{v_{1},\ldots,v_{\max\{s,t\}+1}\}$ denote the excluded vertices in $G_{n+1}$ for each of the $\max\{s,t\}+1$ subgraphs in $\mathcal{R}(s,t,n)$.
Assume there exists a subgraph $G^{\prime}$ of $G_{n+1}$ of size $s$ or $t$. If $G^{\prime}$ excludes any $v_{i}\in S$, then $G^{\prime}$ is a subgraph of the $i$th subgraph in $\mathcal{R}(s,t,n)$. Therefore $G^{\prime}$ cannot be a clique of size $s$ or an independent set of size $t$. If $G^{\prime}$ includes all $v_{i}\in S$, then it is at least of size $\max\{s,t\}+1>\max\{s,t\}$ which is a contradiction. ∎
3 Counterexample Checking
Algorithm 1 checks the belonging of a candidate graph $G_{n+1}$ to counterexamples $\mathcal{R}(s,t,n+1)$.
Algorithm 1 Counterexample Checking
  for $\max\{s,t\}+1$ subgraphs $G_{n}$ of size $n$ of graph $G_{n+1}$ do
     if $G_{n}$ is not isomorphic to some $G^{\prime}_{n}\in\mathcal{R}(s,t,n)$ then
        return false
     end if
  end for
  return true
As opposed to prior algorithms exponentially related to $s$ and $t$, this algorithm requires a linear number of checks. However, other algorithms require no isomorphism checking. This algorithm requires a number of isomorphism calls linearly related to $s$ and $t$, and the isomorphism must also be found to exist or not exist from a set of $k$ possible candidates, giving an upper bound of $k(\max\{s,t\}+1)$ checks in the naive implementation.
Additionally, this algorithm requires knowledge of $\mathcal{R}(s,t,n)$. If some subset of $\mathcal{R}(s,t,n)$ is used, or the set used is not known to be exhaustive, a false return value is possible for a candidate $G_{n+1}\in\mathcal{R}(s,t,n+1)$.
4 One-Vertex Extension
Algorithm 2 extends from $\mathcal{R}(s,t,n)$ to $\mathcal{R}(s,t,n+1)$.
Algorithm 2 One-Vertex Extension
  for all $G_{n}\in\mathcal{R}(s,t,n)$ do
     Define $G_{n+1}:=G_{n}+\{v_{n+1}\}$
     for all $i\in[1,n]$ do
        Graph $G_{n-1}:=G_{n}-v_{i}$
        for all $G^{\prime}_{n}\in\mathcal{R}(s,t,n)$ s.t $G^{\prime}_{n}\neq G_{n},G_{n-1}\subset G^{\prime}_{n}$ do
           Assign edges between $G_{n-1}$ and $v_{n+1}$ s.t. $G_{n-1}+\{v_{n+1}\}=G^{\prime}_{n}$
           Add $G_{n+1}$ to $\mathcal{R}(s,t,n+1)$ if it is a counterexample.
           Connect $v_{n+1}$ to $v_{i}$.
           Add this new $G_{n+1}$ to $\mathcal{R}(s,t,n+1)$ if it is a counterexample.
        end for
     end for
  end for
  return $\mathcal{R}(s,t,n+1)$
This algorithm iterates over at most $2k^{2}n$ candidates. Other extension algorithms look at all ways to connect a new vertex ($2^{n}$) to an existing counterexample, meaning they search $k2^{n}$ candidates. Both extension algorithms in conjunction allow for the improved choice of $\min\{2k^{2}n,k2^{n}\}$ candidates. Additionally, this algorithm explores all possible graphs with at least 2 subgraphs in $\mathcal{R}(s,t,n)$. Because $\mathcal{R}(s,t,n+1)$ is the set of all graphs with $\max\{s,t\}+1$ subgraphs in $\mathcal{R}(s,t,n)$, this exhaustively find all graphs in $\mathcal{R}(s,t,n+1)$.
Regarding checking counterexamples, existing algorithms or Algorithm 1 can be used. When using Algorithm 1, only $\max\{s,t\}-1$ additional iterations are needed as two subgraphs have already been verified.
Running Algorithm 2 on a subset of $\mathcal{R}(s,t,n)$, only graphs with at least $\max\{s,t\}+1$ subgraphs in the subset are generated when using Algorithm 1 to check, and at least 2 subgraphs in the subset when checking with other schemes. This is a stricter limitation than other extension algorithms that generate graphs with at least 1 subgraph in the subset.
An additional cost of this algorithm is that, in my implementation, the condition of finding that $G_{n-1}\subset G^{\prime}_{n}$ requires calls to an isomorphism checker.
5 Results
I wrote a single-threaded python implementation of the high level algorithms provided in Sections 3 and 4. Code was ran on a laptop with a single intel core i7. The code, available at https://github.com/aLehav/ramsey-ove-and-check, is open-sourced and fully documented. Algorithms used and details regarding the code are in Section A. I also decremented to subsets of $\mathcal{R}(s,t,n-1)$ from given subsets of some $\mathcal{R}(s,t,n)$ to broaden the search space, as done by McKay, Exoo, and Radziszowski [5, 6].
5.1 R(4,6,36)
There are no counterexamples of size 36 with at least 7 subgraphs strictly within the 37 known $\mathcal{R}(4,6,35)$ examples [7]. This took 20 seconds to verify.
5.2 R(5,5,43)
There are no counterexamples of size 43 with at least 6 subgraphs strictly within the 656 known $\mathcal{R}(5,5,42)$ examples[7]. This took 9 minutes and 5 seconds to verify.
5.3 Decrementing
I decremented the subset of $\mathcal{R}(4,6,35)$ to a subset of $\mathcal{R}(4,6,34)$ and the subset of $\mathcal{R}(5,5,42)$ to a subset of $\mathcal{R}(5,5,41)$ and tried reconstructing to find new counterexamples. None were found.
6 Conclusion
I introduce a new theorem and way to look at Ramsey theory counterexamples with regard to smaller counterexamples. I provide algorithms that leverage this, and implement them to verify known results. This new approach may assist in novel methods for one-vertex extension or gluing that yield the complete sets $\mathcal{R}(4,6,36)$, $\mathcal{R}(5,5,43)$, or $\mathcal{R}(3,10,40)$ in the future.
Appendix A Code Implementation
Much of the work leveraged the NetworkX library, in particular the VF2++ algorithm [8, 9]. The exact libraries and data structures used may affect the runtime of these algorithms.
A.1 Algorithms
There are three bottlenecks in the practical implementation of Algorithm 1 and Algorithm 2.
  1. 1.
    Algorithm 1 determines if a candidate graph has $\max\{s,t\}+1$ subgraphs of size $n$ that are isomorphic to any one of $k$ counterexamples.
  2. 2.
    Algorithm 2 finds all $k$ counterexamples sharing a subgraph of size $n-1$ with a given candidate graph.
  3. 3.
    In Algorithm 2, after finding a counterexample sharing a subgraph of size $n-1$, the algorithm connects a vertex to recreate an isomorphism to this counterexample and checks if the new $G_{n+1}$ is a counterexample.
Algorithm 1 must determine if a candidate graph has $\max\{s,t\}+1$ subgraphs that are isomorphic to a counterexample. As mentioned, the naive implementation may iterate each of $\max\{s,t\}+1$ arbitrarily chosen subgraphs of $G_{n+1}$ over each of the $k$ counterexamples to check if they are isomorphic. However, by pre-computing the sorted edge degree of each of the $k$ counterexamples, time is saved by only checking a candidate subgraph against the subset of counterexamples with a matching sorted edge degree. This hashing significantly speeds up the code. Using the count of triangles through each vertex or number of all subgraphs of size 3 through a given vertex lowers the number of counterexamples to check against to 1 or 2 and only adds a runtime component of $O(n^{3})$ per candidate.
Additionally, if for some $i\in[1,n+1]$, $G_{n+1}-\{v_{i}\}$ is in $\mathcal{R}(s,t,n)$, then any subgraph of $G_{n+1}$ of size $n$ must have a subgraph of size $n-1$ in $\mathcal{R}(s,t,n-1)$. This is because all subgraphs of size $n$ must have an overlapping subgraph of size $n-1$ with $G_{n+1}-\{v_{i}\}$. With this knowledge, we compose a mapping $\Psi$ from graphs in $\mathcal{R}(s,t,n-1)$ to sets of neighbors of the $n$th vertex that would produce a graph in $\mathcal{R}(s,t,n)$. This is done via Algorithm 5, described later.
Algorithm 2 could then work by iterating over every graph $G_{n}$ in $\mathcal{R}(s,t,n)$ and every index $i\in[1,n]$. For each set of neighbors in $\Psi(G_{n}-\{v_{i}\})$, $G_{n}-\{v_{i}\}$ can be connected to $v_{n+1}$ in an isomorphically equivalent way. Now, $G_{n+1}$ has two subgraphs verified to be in $\mathcal{R}(s,t,n)$. If for $\max{s,t}-1$ additional indices $j\notin\{i,n+1\}$, $\Psi(G_{n+1}-\{v_{n+1},v_{j}\})$ includes the neighbors of $v_{n+1}$ in $G_{n+1}-\{v_{n+1}\}$, then $G_{n+1}-\{v_{j}\}\in\mathcal{R}(s,t,n)$ and so $G_{n+1}$ is a counterexample.
This approach significantly reduces the number of counterexamples needed to find isomorphisms to with candidate graphs in both checking and extension. With this approach I developed the more detailed Algorithm 3 for one-vertex extension and Algorithm 4 for checking within that context.
Algorithm 3 Detailed One-Vertex Extension
  Generate mapping $\Psi$
  for all $G_{n}\in R(s,t,n)$, excluding complements when $s=t$ do
     Graph $G_{n+1}:=G_{n}+\{v_{n+1}\}$
     for all $i\in[1,n]$ do
        Index $j:=1+(i+1\>\text{mod}\>n+1)$
        Graph $G_{n-1}:=G_{n}-\{v_{i}\}$
        Graph $G^{\prime}_{n-1}:=G_{n}-\{v_{j}\}$
        Generate Isomorphism $\Phi$ from $G_{n-1}$ to some graph in $\Psi$’s keys
        Generate Isomorphism $\Phi^{\prime}$ from $G^{\prime}_{n-1}$ to some graph in $\Psi$’s keys
        for all Sets of neighbors $M\in\Psi(\Phi(G_{n-1}))$ do
           Create edge $(v_{n+1},v_{k})$ in $G_{n+1}$ for all $v_{k}\in\Phi^{-1}(M)$
           $P:=v_{n+1}$’s neighbors, currently equal to $\Phi^{-1}(M)$
           if $\Phi^{\prime}(P\text{ excluding $v_{j}$})\in\Psi(\Phi^{\prime}(G^{\prime}_{n-1}))$ then
              Call Algorithm 4 with all local variables
           end if
           Add edge $(v_{n+1},v_{i})$ to $G_{n+1}$ and $v_{i}$ to $P$
           if $\Phi^{\prime}(P\text{ excluding $v_{j}$ })\in\Psi(\Phi^{\prime}(G^{\prime}_{n-1}))$ then
              Call Algorithm 4 with all local variables
           end if
        end for
     end for
  end for
  return all found counterexamples
As an aside, $\max_{G_{n-1}}\{|\Psi(G_{n-1}|\}\leqslant k$. It is usually far less than $k$, as in practice this requires all graphs in $\mathcal{R}(s,t,n)$ to share a subgraph of size $n-1$.
Algorithm 4 Detailed Counterexample Checking
  Tested subgraph count $t:=2$
  Index $k:=j$
  while $t\leqslant\max\{s,t\}$ do
     $k=1+(k+1\>\text{mod}\>n+1)$
     Graph $G^{*}_{n-1}:=G_{n}-\{v_{k}\}$
     Generate Isomorphism $\Phi^{*}$ from $G^{*}_{n-1}$ to some graph in $\Psi$’s keys
     if $\Phi^{*}(P\text{ excluding $k$})\notin\Psi(\Phi^{*}(G^{*}_{n-1}))$ then
        $G_{n+1}$ is not a counterexample
     end if
  end while
  $G_{n+1}$ is a counterexample, add it to $\mathcal{R}(s,t,n+1)$
Algorithm 5 generates $\Psi$. The reason that this algorithm must find all isomorphisms whereas the above algorithms do not is in cases where a vertex is connected via one mapping of a given counterexample but checked from another. If only one isomorphism is added to the mapping’s output sets, a valid counterexample graph may incorrectly be labeled as a false negative.
Algorithm 5 Mapping construction
  Generate blank mapping $\Psi$
  for all $G_{n}\in\mathcal{R}(s,t,n)$ do
     for all $i\in[1,n]$ do
        $G_{n-1}:=G_{n}-\{v_{i}\}$
        $M_{i}:=v_{i}$’s neighbors
        if $\Psi$’s domain contains a graph $G^{\prime}_{n-1}$ isomorphic to $G_{n-1}$ then
           for all Isomorphisms $\Phi$ from $G_{n-1}$ to $G^{\prime}_{n-1}$ do
              Add $\Phi(M_{i})$ to $\Psi(\Phi(G_{n-1}))$
           end for
        else
           for all Automorphisms $\Phi$ from $G_{n-1}$ to itself do
              Add $\Phi(M_{i})$ to $\Psi(\Phi(G_{n-1}))$
           end for
        end if
     end for
  end for
  return $\Psi$
A.2 Runtime
The vast majority of the runtime of this algorithm is spent on finding isomorphisms as well as determining their existence. As noted in Section A.1, hashing schemes can be used to minimize the number of counterexamples a candidate must compare against. In practice, hashing based on sorted vertex degree lowers the maximum number of valid counterexamples to 17 for $R(4,6,41)$ and 86 for $R(5,5,41)$. Hashing based on the number of triangles per vertex runs in equivalent time with NetworkX [8] and decreases this to at most 2 valid counterexamples for both sets. Hashing based on all subgraphs of size 3 through each vertex takes measurably longer but reduces it down to 1 valid counterexample to check against for both sets. An ideal hashing algorithm would not only generate 1 valid counterexample, but in doing so also begin the isomorphic assignment. This is an important direction of future work.
Triangle counts for a graph of size $n$ runs in $O(n^{3})$. I use VF2++ [9] for isomorphism, making my time to hash and find the isomorphism of a single graph $h:=O(n^{3}+1\cdot VF2++)$. This gives, for an arbitrary hashing scheme, $h=O(s+(C+N-1)\cdot\Phi)$ when the scheme runs in $s$ time and leaves at most $C$ valid counterexamples when finding $N$ isomorphisms from the candidate to counterexample and isomorphism generation takes $\Phi$ time for an arbitrary scheme. For dictionary construction, I aim to find the maximum $N$ per graph, and so denote it $h^{\prime}$ in the following analysis.
If I extend from $\mathcal{R}(s,t,n)$ of size k, it takes $O(kh^{\prime})$ to generate $\Psi$, iterate over $2kn\max_{G_{n-1}}\{|\Psi(G_{n-1}|\}=O(k^{2}n)$ candidate graphs, and for each take $O(\max\{s,t\}h)$ time to assign edges and check if a counterexample, for an overall runtime of $O(k^{2}n\max\{s,t\}h+kh^{\prime})$.
Acknowledgments
Thanks to Afia Anjum for her endless patience in listening to me talk about these algorithms.